11 less than 6 times an unknown is equal to -5, what is the unknown?

Answer by Cromlix(4381)

Hi there,
11 less than 6 times an unknown is equal to -5
Make the unknown = 'x'
6x - 11 = -5
6x = -5 + 11
6x = 6
x = 6/6
x = 1.
Unknown = 1
Hope this helps :-)